OpenStudy (anonymous):

write the slope intercept equation for the line with slope 3/4 and y intercept (0,-5)

OpenStudy (katrinakaif):

plug the given slope as "m" in the equation: (y-y1) = m(x-x1) and (x1,y1) = (0, -5) Solve =)

OpenStudy (anonymous):

y=3/4x-5